# Random Transverse Field Effects on Magnetic Noise in Spin Systems

### v1

@@ IV.2 Results
Figure 5: Simulation on the three-dimensional CSI and RTFSI models: (a) Top: Unit cell of the pyrochlore lattice, consisting of 16 spin sites and 4 non-touching tetrahedra (red). The directions of the four easy axes are also shown. Bottom: Schematic plot of a hopping monopole in CSI. The middle tetrahedron hosts a monopole (1-in-3-out configuration); while hopping to the right is permitted without increasing the energy, hopping to the left creates a double monopole (all-in configuration) and is hence energetically suppressed at low temperature. (b) $\mathrm{PSD}(\omega)$ with varying values of $h$ for RTFSI, $r=0.08$ and $T=0.3$. The PSD for the SM dynamics of the classical model is also shown for reference. Inset: $\omega^{2}\,\mathrm{PSD}(\omega)$, focusing on the intermediate region and scaled by a factor of $10^{7}$. (c) Values of $A_{\infty}/A_{0}$, obtained by fitting the two regions of the PSD bounded by the dashed vertical lines to the form $\mathrm{PSD}(\omega)=A\omega^{-2}$, as a function of $h$; results for $r=0.02$ (red), $r=0.04$ (gray), $r=0.06$ (blue), $r=0.08$ (purple), $T=0.3$ (solid lines, filled markers) and $T=0.35$ (dotted lines, empty markers) are shown. The dashed horizontal line represents the value $A_{\infty}/A_{0}\approx 1.53$ measured for CSI, see main text.
@@ V.1 Setup
$\mathcal{H}_{\mathrm{RTFSI}}=\sum_{\langle ij\rangle}S_{i}^{z}S_{j}^{z}+\sum_{\{\alpha\}}S_{\alpha}^{x}\,,$ (24)
@@
We probe the dynamics of the RTFSI model using the same modified Monte-Carlo method outlined at the end of Sec. II. Namely, we sample the Master equation Eq. (10). Without the random transverse fields, such dynamics would be identical to the ‘standard model’ (SM) dynamics of the CSI [55, 52]. This is the starting point of our analysis. For simplicity, a strong magnetic anisotropy is assumed also for the spins affected by the transverse fields, and only their easy-axis ($z$) component is taken to contribute to the total magnetization of the system. The RTFSI system is isotropic in the $x$-, $y$- and $z$-directions; for concreteness, we choose to compute the PSD using the average over $m_{x}$, $m_{y}$ and $m_{z}$, or equivalently the magnetization along the [111] direction. See App. F for simulation details.
@@ V.2 Results
We start by discussing the behavior of the SM dynamics. Similarly to the KIC, at low temperature the magnetic noise of the CSI in equilibrium is dominated by the contribution from the random walking point-like excitations known as magnetic monopoles. However, such a random walk is constrained: for a given isolated monopole, the move corresponding to flipping the minority spin of the corresponding 1-in-3-out (3-in-1-out) tetrahedron is energetically suppressed (see bottom panel of Fig. 5a), as it creates a double-monopole corresponding to an all-out (all-in) tetrahedron. This is different from the 1D scenario where the domain walls can freely move in all directions.
@@
We now turn to discuss the results for the RTFSI model, presented in Figs. 5b and 5c. Similarly to the 1D case, the introduction of the random transverse fields increases the deviation between the low and high frequency regimes of the PSD, and hence enlarges $A_{\infty}/A_{0}$ with respect to the classical behavior. Fig. 5c reveals this quantity to increase as a function of $h$ and $r$ and decrease with $T$, also in agreement with our 1D results (see Fig. 3).
@@ Acknowledgements.
We would like to thank John Chalker and Orazio Scarlatella for useful discussions. This work was supported in part by the Engineering and Physical Sciences Research Council (EPSRC) grants No. EP/T028580/1 and No. EP/V062654/1 (C.C.). This work was in part supported by the Deutsche Forschungsgemeinschaft under Grants No. SFB 1143 (Project No. 247310070) and the cluster of excellence ct.qmat (EXC 2147, Project No. 390858490). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ A.2 Heat capacity
For $n\geq 2$, $\bar{H}^{(n)}$ is dual to $H^{(n-1)}$ in Eq. (5) and they share the same spectrum. On the other hand, $\bar{H}^{(1)}=-\bar{S}_{1}^{z}$ is a simple two-level system, with the excited state signifying an anti-aligned pair of nearest-neighbor spins in the original system, i.e., a domain wall. (At low temperatures, the thermal occupancy of the excited states is determined by the corresponding Boltzmann factor $e^{-2\beta}$; in this way, one can then straightforwardly verify the claim about the excitation density and typical separation in the KIC made in Sec. II.1.)
@@ A.4 Discussion
In the KIC, the PSD of the system per spin takes the form [64, 65, 66, 52]
@@ C.1 Model
To facilitate discussion, we define $s(i)=n$ if site $i$ on the RW-lattice represents an $n$-cluster, and $s(i)=0$ if it represents a classical site, and define $\tilde{W}_{s(i),s(j)}\equiv W_{ij}\delta_{i,j\pm 1}$. Using Eq. (12) in the main text, we have
@@
As mentioned in the main text, we also need to adjust the step lengths between adjacent sites in the RW-lattice to account for the (spatial and magnetic) size of the quantum clusters. We define $a_{ij}$ to be the step length between sites $i$ and $j$, representing the difference in magnetization induced via transitions between these sites ($a_{ij}=a_{ji}$). We define $\tilde{a}_{s(i),s(j)}=a_{ij}\delta_{i,j\pm 1}$ as:
@@ C.2 Diffusion constants
being the average step length. Eq. (38) suggests that the value of the diffusion constant $D_{\infty}$ is given by the product of $\bar{a}^{2}$ times the same quantity defined for a random walk with the same Master equation Eq. (16) but happening on a standard lattice (i.e., with uniform unit bond length). The latter quantity can then be calculated using Eqs. (46) to (50) in Ref. 71 and the relationship $\tilde{W}_{n,n^{\prime}}\tilde{P}_{n}^{0}=\tilde{W}_{n^{\prime},n}\tilde{P}^{0}_{n^{\prime}}$. Altogether, we obtain
@@
These analytical results are used in plotting the inset of Fig. 2b as well as the results for $D_{0}/D_{\infty}$ in Fig. 3. There, evaluations of the sums over sites in Eqs. (37), (39) and (40) are performed on a RW-lattice corresponding to a $10^{6}$-spin system.
@@ Appendix E Magnetization dynamics in CSI
Figure 10: Magnetization dynamics in CSI: (a) Results for the free, constrained and monopolar random walks (see main text) on the diamond lattice. The free walk behaves diffusively at all times, while the constrained and monopolar random walks are diffusive at short and long times, with a crossover between the two distinct diffusion constants at intermediate times. The three horizontal black lines represent the different diffusive behaviors observed in these models, with $M(t)/t=3$ (dashed), $2.4$ (dotted) and $1.85$ (dash-dotted). (b) and (c) The value of $A_{\infty}/A_{0}$ obtained by fitting the low- and high-frequency regions of the PSD (see Sec. V and Fig. 5b), simulated with system sizes $6\leq L\leq 18$ and temperatures $0.2\leq T\leq 0.4$.
@@
To understand the effect of such constraints, we simulated and compared the following random walks on the diamond lattice: (i) a free, standard random walk on the diamond lattice with unit bond length, with the transition rates across each bonds set to unity, hereafter referred to as the ‘free walk’; (ii) a constrained random walk, where the walker is always allowed to backtrack but finds one of the three other directions to be blocked at random (and chosen anew each time, at the time of hopping of the walker), hereafter referred to as the ‘constrained walk’; (iii) the monopolar random walk happening in CSI in the single-excitation limit, hereafter referred to as the ‘standard model’ (SM) random walk. The last walk is simulated following the procedure described in Ref. 57. We randomly initialize a spin system with $40\times 40\times 40$ unit cells ($10^{6}$ spins; see App. F); after that, perform single-spin flip Monte-Carlo updates at $T=0$ until no monopoles are present in the system, giving us a ground state configuration. Flipping a spin selected at random creates a monopole-anti-monopole pair; by fixing one of them and allowing the other to move in the system, we are able to calculate the MSD $M(t)$, defined at the beginning of Sec. III.2 in the main text, from the magnetization change of the system incurred by such a walk.
Fig. 10a summarizes the data obtained for the ratio $M(t)/t$. For the free walk, one has trivially $M(t)/t=4$ from the coordination number of the diamond lattice. For both the constrained walk and the SM, a difference between the short- and long-time diffusive behavior appears. Both exhibit $D_{0}\equiv\lim_{t\to 0}M(t)/t=3$, attributed to the 3 out of 4 available hopping directions the walker can take in its first step. As for $D_{\infty}\equiv\lim_{t\to\infty}M(t)/t$, the results obtained from our simulations point to $D_{\infty}=2.4$ for the constrained walk [45], and $D_{\infty}\approx 1.85$ for the SM. In the latter case, however, the presence of finite size effects at $M(t)\gtrsim L^{2}$ makes the accurate extraction of $D_{\infty}$ difficult; hence, we estimate a $1.85\leq D_{\infty}\leq 1.90$ for SM.
The constrained walk can be exactly solved using the methods developed in Chapter 4 of Ref. 45. In particular, it is a random walk with correlations over two successive steps (as the probability of a transition depends on the previous one — recall that backtracking is always possible). Eq. 4.25 in the reference gives the exact solution of the MSD of such random walks and resembles Eq. (17) in the main text:
$M(t)=2.4t+0.18(\mathrm{e}^{-t/\tau}-1)\,,$ (43)
@@
The SM random walk differs from the constrained walk by the important aspect of spin correlations [22]. In the ground state, they imply correlated blocked directions for the monopolar walker. Moreover, the Dirac string — a chain of flipped spins along the path — left by the mobile monopole also implies a memory effect of the random walk dynamics, which grows with time as the monopole leaves behind an increasingly longer trail. As can be seen from Fig. 10a, comparing with the constrained walk, such effects seem to come in at around $t\sim 10^{0}$, preventing the walker from exploring farther sites (hence reducing $M(t)/t$) and leaving a long relaxation tail in the MSD of the SM before it converges to the long-time asymptotic diffusive behavior.
The above random walk analysis does not take into account the presence of finite temperature. Moreover, one may worry about the relatively small system sizes accessible in the spin simulation results discussed in the main text ($L<20$, comparing with $L=40$ used in the above SM random walk simulations and $L\sim 10^{3}$ in 1D; see App. F). For this reason, in Fig. 10b and 10c, we investigate the dependence of the shape of the PSD obtained from the spin simulations in the main text with temperature $T$ and system size $L$, by comparing the values $A_{\infty}/A_{0}$ using the same method and fitting range as in the main text (see Sec. V.2 and Fig. 5b). As can be seen from the figures, $A_{\infty}/A_{0}$ has little variation within the system sizes considered and the temperature range, so long as $T\leq 0.35$. The upturn at higher temperatures can be attributed to a decreasing magnetic relaxation time which causes the low-frequency region to deviate from the $\mathrm{PSD}\propto\omega^{-2}$ behavior — see the discussion in Sec. A.4.
@@ F.1.1 One-dimensional RTFIC
To obtain the results presented in Fig. 1, we performed Monte-Carlo simulations as described at the end of Sec. II. In the simulations, we consider $N$ spin-flip attempts as 1 MC-iteration, which is then associated with the passage of 1 unit of MC time. We simulate a system with $2000$ spins and periodic boundary conditions, initially in a random state (with classical spins randomly aligned and the quantum clusters in a random eigenstate of their respective Hamiltonians). A cutoff cluster size $n=8$ is used, which is sensible in our range for $r$ (the probability of finding clusters of size $n>8$ in a $2000$-spin system with $r=0.2$, for example, is $\sim 0.01$). After $2\times 10^{4}$ equilibrium iterations, we run the system for an extra $2^{20}\approx 10^{6}$ iterations, measuring the magnetization $m(t)$ every $1/50$ steps (i.e., every $40$ spin-flip attempts). We then employ the scipy.signal.welch function [72] to compute the PSDs presented in Fig. 1a using $5242882^{19}=$ 524288 periodograms. Each PSD curve is averaged over $12$ MC simulations with different realizations of the random transverse field.

### v2

@@ IV.2 Results
Figure 5: Simulation on the three-dimensional CSI and RTFSI models: (a) Top: Unit cell of the pyrochlore lattice, consisting of 16 spin sites and 4 non-touching tetrahedra (red). The directions of the four easy axes are also shown. Bottom: Schematic plot of a hopping monopole in CSI. The middle tetrahedron hosts a monopole (1-in-3-out configuration); while hopping to the right is permitted without increasing the energy, hopping to the left creates a double monopole (all-in configuration) and is hence energetically suppressed at low temperature. (b) $\mathrm{PSD}(\omega)$ with varying values of $h$ for RTFSI, $r=0.08$ and $T=0.3$. The PSD for the standard model dynamics of the classical model is also shown for reference black dash-dot line. Inset: $\omega^{2}\,\mathrm{PSD}(\omega)$, focusing on the intermediate region and scaled by a factor of $10^{7}$. (c) Values of $A_{\infty}/A_{0}$, obtained by fitting the two regions of the PSD bounded by the dashed vertical lines to the form $\mathrm{PSD}(\omega)=A\omega^{-2}$, as a function of $h$; results for $r=0.02$ (red), $r=0.04$ (gray), $r=0.06$ (blue), $r=0.08$ (purple), $T=0.3$ (solid lines, filled markers) and $T=0.35$ (dotted lines, empty markers) are shown. The dashed horizontal line represents the value $A_{\infty}/A_{0}\approx 1.53$ measured for CSI, see main text.
@@ V.1 Setup
$\mathcal{H}_{\mathrm{RTFSI}}=\sum_{\langle ij\rangle}S_{i}^{z}S_{j}^{z}{-h}\sum_{\{\alpha\}}S_{\alpha}^{x}\,,$ (22)
@@
We probe the dynamics of the RTFSI model using the same modified Monte-Carlo method outlined at the end of Sec. II. Namely, we sample the Master equation Eq. (8). Without the random transverse fields, such dynamics would be identical to the ‘standard model’ dynamics of the CSI [58, 54]. This is the starting point of our analysis. For simplicity, a strong magnetic anisotropy is assumed also for the spins affected by the transverse fields, and only their easy-axis ($z$) component is taken to contribute to the total magnetization of the system. The RTFSI system is isotropic in the $x$-, $y$- and $z$-directions; for concreteness, we choose to compute the PSD using the average over $m_{x}$, $m_{y}$ and $m_{z}$, or equivalently the magnetization along the [111] direction. See App. F for simulation details.
@@ V.2 Results
We start by discussing the behavior of the standard model dynamics. Similarly to the classical kinetic Ising chain, at low temperature the magnetic noise of the CSI in equilibrium is dominated by the contribution from the random walking point-like excitations known as magnetic monopoles. However, such a random walk is constrained: for a given isolated monopole, the move corresponding to flipping the minority spin of the corresponding 1-in-3-out (3-in-1-out) tetrahedron is energetically suppressed (see bottom panel of Fig. 5a), as it creates a double-monopole corresponding to an all-out (all-in) tetrahedron. This is different from the 1D scenario where the domain walls can freely move in all directions.
@@
We now turn to discuss the results for the RTFSI model. Here, we choose to simulate values of $r$ much less than those used in 1D. This is due to the fact that the effects of the quantum clusters are more prominent in the three dimensional spin ice model than in 1D, owing to the larger coordination number (which both leads to larger clusters and a larger number of neighboring classical spins – an isolated quantum spin, for example, has two neighbors in 1D but six in 3D). Moreover, the choice also helps contain the computational efforts required for our study. Specifically, we restrict to $0<r\leq 0.08$ and we study a broader range of $h$ so as to better resolve the effects of the quantum clusters. The results are presented in Figs. 5b and 5c; similarly to the 1D case, the introduction of the random transverse fields increases the deviation between the low and high frequency regimes of the PSD, and hence enlarges $A_{\infty}/A_{0}$ with respect to the classical behavior. Fig. 5c reveals this quantity to increase as a function of $h$ and $r$ and decrease with $T$, also in agreement with our 1D results (see Fig. 3).
@@ VI Discussion and outlook
Data Availability Statement
All data produced and presented in this work were obtained by well-known methods that have been described in detail in the main text and App. F, and can be readily reproduced. They can also be made available upon reasonable request.
@@ Acknowledgements.
We would like to thank John Chalker, Orazio Scarlatella, Jonathan N. Hallén and Nilotpal Chakraborty for useful discussions. This work was supported in part by the Engineering and Physical Sciences Research Council (EPSRC) grants No. EP/T028580/1 and No. EP/V062654/1 (C.C.). This work was in part supported by the Deutsche Forschungsgemeinschaft under Grants No. SFB 1143 (Project No. 247310070) and the cluster of excellence ct.qmat (EXC 2147, Project No. 390858490). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ A.2 Heat capacity
For $n\geq 2$, $\bar{H}^{(n)}$ is dual to $H^{(n-1)}$ in Eq. (3) and they share the same spectrum. On the other hand, $\bar{H}^{(1)}=-\bar{S}_{1}^{z}$ is a simple two-level system, with the excited state signifying an anti-aligned pair of nearest-neighbor spins in the original system, i.e., a domain wall. (At low temperatures, the thermal occupancy of the excited states is determined by the corresponding Boltzmann factor $e^{-2\beta}$; in this way, one can then straightforwardly verify the claim about the excitation density and typical separation in the classical Ising chain made in Sec. II.1.)
@@ A.4 Discussion
In the kinetic classical Ising chain, the PSD of the system per spin takes the form [67, 68, 69, 54]
@@ C.1 Model
To facilitate discussion, we define $s(i)=n$ if site $i$ on the random walk lattice represents an $n$-cluster, and $s(i)=0$ if it represents a classical site, and define $\tilde{W}_{s(i),s(j)}\equiv W_{ij}\delta_{i,j\pm 1}$. Using Eq. (10) in the main text, we have
@@
As mentioned in the main text, we also need to adjust the step lengths between adjacent sites in the random walk lattice to account for the (spatial and magnetic) size of the quantum clusters. We define $a_{ij}$ to be the step length between sites $i$ and $j$, representing the difference in magnetization induced via transitions between these sites ($a_{ij}=a_{ji}$). We define $\tilde{a}_{s(i),s(j)}=a_{ij}\delta_{i,j\pm 1}$ as:
@@ C.2 Diffusion constants
being the average step length. Eq. (36) suggests that the value of the diffusion constant $D_{\infty}$ is given by the product of $\bar{a}^{2}$ times the same quantity defined for a random walk with the same Master equation Eq. (14) but happening on a standard lattice (i.e., with uniform unit bond length). The latter quantity can then be calculated using Eqs. (46) to (50) in Ref. 74 and the relationship $\tilde{W}_{n,n^{\prime}}\tilde{P}_{n}^{0}=\tilde{W}_{n^{\prime},n}\tilde{P}^{0}_{n^{\prime}}$. Altogether, we obtain
@@
These analytical results are used in plotting the inset of Fig. 2b as well as the results for $D_{0}/D_{\infty}$ in Fig. 3. There, evaluations of the sums over sites in Eqs. (35), (37) and (38) are performed on a random walk lattice corresponding to a $10^{6}$-spin system.
@@ Appendix E Magnetization dynamics in CSI
Figure 10: Magnetization dynamics in CSI: (a) Results for the free, constrained and standard model random walks (see main text) on the diamond lattice. The free walk behaves diffusively at all times, while the constrained and standard model random walks are diffusive at short and long times, with a crossover between the two distinct diffusion constants at intermediate times. The three horizontal black lines represent the different diffusive behaviors observed in these models, with $M(t)/t=3$ (dashed), $2.4$ (dotted) and $1.85$ (dash-dotted). (b) and (c) The value of $A_{\infty}/A_{0}$ obtained by fitting the low- and high-frequency regions of the PSD (see Sec. V and Fig. 5b), simulated with system sizes $6\leq L\leq 18$ and temperatures $0.2\leq T\leq 0.4$.
@@
To understand the effect of such constraints, we simulated and compared the following random walks on the diamond lattice: (i) a free, standard random walk on the diamond lattice with unit bond length, with the transition rates across each bonds set to unity, hereafter referred to as the ‘free walk’; (ii) a constrained random walk, where the walker is always allowed to backtrack but finds one of the three other directions to be blocked at random (and chosen anew each time, at the time of hopping of the walker), hereafter referred to as the ‘constrained walk’; (iii) the monopolar random walk happening in CSI in the single-excitation limit, hereafter referred to as the ‘standard model’ random walk. The last walk is simulated following the procedure described in Ref. 60. We randomly initialize a spin system with $40\times 40\times 40$ unit cells ($10^{6}$ spins; see App. F); after that, perform single-spin flip Monte-Carlo updates at $T=0$ until no monopoles are present in the system, giving us a ground state configuration. Flipping a spin selected at random creates a monopole-anti-monopole pair; by fixing one of them and allowing the other to move in the system, we are able to calculate the MSD $M(t)$, defined at the beginning of Sec. III.2 in the main text, from the magnetization change of the system incurred by such a walk.
Fig. 10a summarizes the data obtained for the ratio $M(t)/t$. For the free walk, one has trivially $M(t)/t=4$ from the coordination number of the diamond lattice. For both the constrained walk and the standard model, a difference between the short- and long-time diffusive behavior appears. Both exhibit $D_{0}\equiv\lim_{t\to 0}M(t)/t=3$, attributed to the 3 out of 4 available hopping directions the walker can take in its first step. As for $D_{\infty}\equiv\lim_{t\to\infty}M(t)/t$, the results obtained from our simulations point to $D_{\infty}=2.4$ for the constrained walk [47], and $D_{\infty}\approx 1.85$ for the standard model. In the latter case, however, the presence of finite size effects at $M(t)\gtrsim L^{2}$ makes the accurate extraction of $D_{\infty}$ difficult; hence, we estimate a $1.85\leq D_{\infty}\leq 1.90$ for standard model.
The constrained walk can be exactly solved using the methods developed in Chapter 4 of Ref. 47. In particular, it is a random walk with correlations over two successive steps (as the probability of a transition depends on the previous one — recall that backtracking is always possible). Eq. 4.25 in the reference gives the exact solution of the MSD of such random walks and resembles Eq. (15) in the main text:
$M(t)=2.4t+0.18(1-\mathrm{e}^{-t/\tau})\,,$ (41)
@@
The standard model random walk differs from the constrained walk by the important aspect of spin correlations [22]. In the ground state, they imply correlated blocked directions for the monopolar walker. Moreover, the Dirac string — a chain of flipped spins along the path — left by the mobile monopole also implies a memory effect of the random walk dynamics, which grows with time as the monopole leaves behind an increasingly longer trail. As can be seen from Fig. 10a, comparing with the constrained walk, such effects seem to come in at around $t\sim 10^{0}$, preventing the walker from exploring farther sites (hence reducing $M(t)/t$) and leaving a long relaxation tail in the MSD of the standard model before it converges to the long-time asymptotic diffusive behavior.
The above random walk analysis does not take into account the presence of finite temperature. Moreover, one may worry about the relatively small system sizes accessible in the spin simulation results discussed in the main text ($L<20$, comparing with $L=40$ used in the above standard model random walk simulations and $L\sim 10^{3}$ in 1D; see App. F). For this reason, in Fig. 10b and 10c, we investigate the dependence of the shape of the PSD obtained from the spin simulations in the main text with temperature $T$ and system size $L$, by comparing the values $A_{\infty}/A_{0}$ using the same method and fitting range as in the main text (see Sec. V.2 and Fig. 5b). As can be seen from the figures, $A_{\infty}/A_{0}$ has little variation within the system sizes considered and the temperature range, so long as $T\leq 0.35$. The upturn at higher temperatures can be attributed to a decreasing magnetic relaxation time which causes the low-frequency region to deviate from the $\mathrm{PSD}\propto\omega^{-2}$ behavior — see the discussion in Sec. A.4.
@@ F.1.1 One-dimensional RTFIC
To obtain the results presented in Fig. 1, we performed Monte-Carlo simulations as described at the end of Sec. II. In the simulations, we consider $N$ spin-flip attempts as 1 MC-iteration, which is then associated with the passage of 1 unit of MC time. We simulate a system with $2000$ spins and periodic boundary conditions, initially in a random state (with classical spins randomly aligned and the quantum clusters in a random eigenstate of their respective Hamiltonians). A cutoff cluster size $n=8$ is used, which is sensible in our range for $r$ (the probability of finding clusters of size $n>8$ in a $2000$-spin system with $r=0.2$, for example, is $\sim 0.01$). After $2\times 10^{4}$ equilibrium iterations, we run the system for an extra $2^{20}\approx 10^{6}$ iterations, measuring the magnetization $m(t)$ every $1/50$ steps (i.e., every $40$ spin-flip attempts). We then employ the scipy.signal.welch function [75] to compute the PSDs presented in Fig. 1a using $5242882^{19}=$ 524288 periodograms. Each PSD curve is averaged over $12$ MC simulations with different realizations of the random transverse field. The results in the limit $h\to 0$, used in the main text to compare with the classical scenario, are obtained by setting $h=0.01$.